OpenStudy (anonymous):

need help solving \[(-4x \sqrt{x})(2x \sqrt{x})\]

OpenStudy (cggurumanjunath):

-4*x*2*x=-8x^2

OpenStudy (cggurumanjunath):

sqrt(x)*sqrt(x)=x

OpenStudy (cggurumanjunath):

-8x^2*x=-8*x^3

OpenStudy (anonymous):

thank you

OpenStudy (cggurumanjunath):

Anytime. :) welcome.